OpenStudy (anonymous):

estimate the quantity using linear approximation 1/sqrt (99)- 1/sqrt(98)

OpenStudy (zarkon):

find the linearization of \(\displaystyle\frac{1}{\sqrt{x}}\)

OpenStudy (anonymous):

im not sure what the linearization would be

OpenStudy (zarkon):

let \[f(x)=\displaystyle\frac{1}{\sqrt{x}}\] then \[L(x)=f'(a)(x-a)+f(a)\]

OpenStudy (anonymous):

ok i got that and then i subtract the two answers right?

OpenStudy (zarkon):

sure

OpenStudy (anonymous):

is -.50000 right ? i am having trouble finding the answer

OpenStudy (zarkon):

no

OpenStudy (zarkon):

i get \[-5\times10^{-4}\]

OpenStudy (anonymous):

oh yea i redid it and got that too thanks